### v2

# Similarity Memory Prior is All You Need for Medical Image Segmentation

Abstract
In recent years, it has been found that “grandmother cells” in the primary visual cortex (V1) of macaques can directly recognize visual input with complex shapes. This inspires us to examine the value of these cells in promoting the research of medical image segmentation. In this paper, we design a Similarity Memory Prior Network (Sim-MPNet) for medical image segmentation. Specifically, we propose a Dynamic Memory Weights-Loss Attention (DMW-LA), which matches and remembers the category features of specific lesions or organs in medical images through the similarity memory prior in the prototype memory bank, thus helping the network to learn subtle texture changes between categories. DMW-LA also dynamically updates the similarity memory prior in reverse through Weight-Loss Dynamic (W-LD) update strategy, effectively assisting the network directly extract category features. In addition, we propose the Double-Similarity Global Internal Enhancement Module (DS-GIM) to deeply explore the internal differences in the feature distribution of input data through cosine similarity and euclidean distance. Extensive experiments on four public datasets show that Sim-MPNet has better segmentation performance than other state-of-the-art methods. Our code is available on https://github.com/vpsg-research/Sim-MPNet.
1 Introduction
Medical image segmentation, as a basic and key technology, is increasingly highlighting its importance in disease diagnosis, treatment planning, and curative effect evaluation [50, 29, 9]. With the rapid development of artificial intelligence and computer vision, constructing automatic and accurate medical image segmentation methods has become a research hotspot [19, 41]. These methods aim to accurately identify and segment the anatomical structures or pathological features of interest from medical images, to provide clinicians with more intuitive and quantitative medical information [43, 21].
Figure 1: Motivation. (a) Schematic diagram of the primate visual system. (b) We use similarity memory priors stored in the prototype memory bank to imitate “grandmother cells” in V1, and utilize these features to directly match various organs in medical images.
Figure 2: Compared with the existing works. CNN passively responds to local texture through convolution kernel. ViT only focuses on high-frequency textures through self-attention. DMW-LA uses similarity memory priors to directly identify and extract category features of organs.
In existing medical image segmentation works, researchers attempt to use Convolutional Neural Network (CNN) or Vision Transformer (ViT) to enable networks to generate images with precise labels [30, 48, 6, 8, 3]. However, although the local receptive field design of CNNs conform to the translational invariance of images, its inductive bias is essentially a passive response to local textures (edges and corners), rather than actively modeling semantic objects (lesions and organs). ViTs require massive data to learn semantic associations of position encoding, while medical image annotation is scarce, leading to model degradation into “pseudo-global modeling” that focuses on high-frequency texture rather than anatomical structure. In addition, CNNs and ViTs belong to data-driven “implicit learning” and lack prior knowledge to quickly identify key features. Recently, researchers have used two-photon calcium imaging technology to discover the presence of curvature detecting cells in the V1 [33, 34]. These curvature detecting cells can be regarded as a type of “grandmother cells” with specific activation conditions, as shown in Figure 7(a). The “grandmother cells” have strong recognition abilities and only respond to specific input stimuli [14, 26, 32]. If the design of medical segmentation networks is not limited to traditional calculation mechanisms and directly focuses on lesions or organs (category features), it may be a promising design paradigm.
Inspired by this, we try to construct the prototype memory bank with different similarity memory priors to simulate the biological mechanism of “grandmother cells”, as shown in Figure 7(b). Specifically, similarity memory priors are matched with organ features in medical images through similarity, and a multi-dimensional mapping relationship based on similarity is constructed to extract rich category features in medical images. Meanwhile, this mapping relationship is strengthened by updating the prototype memory bank. In this process, the prototype memory bank is equivalent to a dynamic prior knowledge repository. With the continuous input of new organ features, the similarity memory priors will be continuously optimized to match organ category features more accurately.
In this paper, we design the Dynamic Memory Weights–Loss Attention (DMW-LA) to directly extract rich category features from medical images. DMW-LA accurately matches the similarity memory priors with the category information in medical images through similarity. This process enables the network to directly capture the abstract category information in medical images, thus significantly enhancing the network’s segmentation ability. In addition, to ensure that the similarity memory priors in the prototype memory bank can accurately identify lesions or organs, we design a novel Weight-Loss Dynamic (W-LD) updating strategy that dynamically updates the prototype memory bank through feature weights and training losses to maintain the effectiveness of similarity memory priors. To deeply explore the hidden patterns and structures in features, we design a Double-Similarity Global Internal Enhancement Module (DS-GIM) to help the network better understand and distinguish subtle differences between different lesions or organs. Finally, we distill the above insights, and construct a Similarity Memory Prior Network (Sim-MPNet) with dual encoder structure by introducing MaxViT [35] to strengthen the learning of global context. In summary, the main contributions of this paper can be summarized as follows:
  • We design a novel Similarity Memory Prior Network (Sim-MPNet), which can fully integrate and utilize category features and global context in medical images, thus improving the accuracy in segmentation tasks.
  • We design Dynamic Memory Weights–Loss Attention (DMW-LA). DMW-LA utilizes prototype memory bank to recognize and extract category information, and dynamically adjusts the semantic cluster features through Weight-Loss Dynamic (W-LD) update strategy. This method is a novel feature extraction paradigm that provides a new perspective for the design of medical image segmentation networks.
  • We design a Double-Similarity Global Internal Enhancement Module (DS-GIM) to help the network better understand the content of medical images. DS-GIM utilizes cosine similarity and euclidean distance to finely distinguish and capture subtle differences between lesions or organs, thus enhancing the network representation ability.
2 Related Work
Medical Image Segmentation. In medical image segmentation, a common method is to segment the image with a full convolutional neural network (FCNN) [15, 11]. Especially after the landmark network U-Net [30] came out, researchers successively launched a series of similar U-shaped Convolutional Neural Network (CNN), such as U-Net++ [48], U-Net3+ [20] and nnU-Net [22]. However, CNNs tend to explore the correlation between local elements, which results in poor performance in segmenting large lesions or organs. Therefore, some researchers use Vision Transformer (ViT) [12] for segmentation, such as MEDT [36], DAE-former [2] and DTMFormer [39]. The self-attention mechanism in ViT focuses on information interaction from a global perspective when processing image information, which lacks exploration of the relationships between local elements. Therefore, some scholars combine CNNs with ViTs to improve the local and global modeling ability of networks, such as TransUNet [8], MixTrans [38], UCTransNet [37] and Transfuse [47]. However, this strategy of integrating local details and global context may be difficult to capture useful category features directly, thus limiting the segmentation performance, as shown in Figure 2. Different from the traditional computing paradigm, our method directly extracts and remembers the representative category information from images, which realizes the accurate segmentation of the target contour.
Prototype Memory Mechanism. The memory bank can capture the cross-time features in network training and has been proven effective in other fields. Santoro et al. [31] used a memory bank to store new training data for few-shot learning, while Yang and Chan [44] employed it to memorize long-term object information for video tracking. Additionally, Caron et al. [7] utilized memory queues to accumulate negative samples for unsupervised learning, and Li et al. [23] proposed Memory Concept Attention (MOCA) to improve the quality of image generation. However, medical images contain various lesions and organs structures, and these structures exhibit significant differences in different individuals and different disease states. For example, the shape, size, position and surrounding tissues of tumors in medical images are ever-changing. The existing prototype memory mechanism lacks sufficient flexibility to update the representation and memory of different structural features in a timely manner. In our work, a novel dynamic memory prior mechanism is designed, which is called DMW-LA. DMW-LA utilizes prototype memory bank to effectively identify the category features of lesions or organs, and dynamically updates the memory bank through the W-LD update strategy. In addition, DS-GIM is also designed to enhance the difference of feature distribution between different lesions or organs.
Figure 3: Overview of Sim-MPNet. The network consists of two encoders, one decoder and one skip connection, and it efficiently models intra-class and global dependencies of input features through the dual encoder structure.
Figure 4: Overview of DMW-LA and W-LD update strategy. (a) DMW-LA consists of intra-cluster interaction and global interaction. (b) The W-LD update strategy adopts weight and loss dual guidance to update the prototype memory bank.
3 Methodology
3.1 Overview
Figure 3 shows the overall structure of the Sim-MPNet. Among them, the encoder 1 consists of four stages DMW-LA. We alternately stack DMW-LA and DS-GIM in the third stage. The encoder 2 consists of a four-stage visual backbone MaxViT [35]. Each encoder uses overlapping convolution operations to achieve down-sampling. The input image will be re-fused with the stage features of each encoder through bilinear interpolation to reduce information loss caused by resolution reduction. After the encoder completes feature extraction, the network fuses the corresponding stage features of the two encoders, as shown in Equation 1.
${E}_{{f}_{i}}={E}_{{f}_{i}}^{1}+{E}_{{f}_{i}}^{2}$ (1)
Where $i=\{1,2,3,4\}$, indicating the encoder stage. ${E}_{{f}_{i}}^{1}$ and ${E}_{{f}_{i}}^{2}$ represent the features of encoder 1 and encoder 2 in stage $i$ respectively. Then, the ${E}_{{f}_{1}}$, ${E}_{{f}_{2}}$ and ${E}_{{f}_{3}}$ is passed into the skip connection, ${E}_{{f}_{4}}$ as the input of the decoder. In skip connection, we connect channel attention (CA) and spatial attention (SA) [40] in series to fuse the corresponding features of the encoder and decoder, as shown in Equation 2.
${O}_{i}=SA(CA({E}_{{f}_{4-i}}+{D}_{{f}_{i}}))$ (2)
Where $i=\{1,2,3\}$, ${D}_{{f}_{i}}$ represents the output of decoder in stage $i$, and ${O}_{i}$ represents the output of skip connection. Then, ${O}_{i}+{D}_{{f}_{i}}$ is used as the input of the next stage. The decoder uses DMW-LA to extract deep semantic information and bilinear interpolation to restore the image to the same resolution as the input image to achieve pixel-level segmentation. Next, we will introduce each module in detail.
3.2 Dynamic Memory Weight-Loss Attention
In this paper, we design Dynamic Memory Weight–Loss Attention (DMW-LA) to extract key category features from medical images, as shown in Figure 4(a). The intra-cluster interaction is responsible for obtaining internal dependencies of category features, while global interaction is responsible for initializing prototype memory bank and modeling global dependencies. Assuming input image ${X}_{1}\in{\mathbb{R}}^{H\times W\times C}$, where $H$, $W$, and $C$ refer to the height, width, and channel dimensions of the image, respectively. First, ${X}_{1}$ is linearly transformed with weights ${W}^{q}$, ${W}^{k}$ and ${W}^{v}$ to generate Query, Key, and Value for global interaction. At the beginning of network training, we perform K-means on the result ${X}^{\prime}$ of global interaction to fill the memory bank. Here, we call the generated semantic clusters as similarity memory prior, and the central point of each cluster is called the similarity core prior. Therefore, the similarity memory prior can be represented as the $\left\{\left\{{E}_{11},{E}_{12},\cdots,{E}_{1c}\right\},\cdots,\left\{{E}_{k1},{E}_{k2},\cdots,{E}_{kc}\right\}\right\}$, and the similarity core prior can be represented as the $\left\{\tilde{{E}_{1}},\tilde{{E}_{2}},\tilde{{E}_{3}},\cdots,\tilde{{E}_{k}}\right\}$. After the initialization of the memory bank is completed, we can supplement prior knowledge for network training. Here, we first calculate the cosine similarity between the Query and the similarity core prior to match the corresponding similarity memory prior. Assuming there is $\left\{{X}^{i}\in{\mathbb{R}}^{N\times C},N\leqslant H\times W\right\}$ in the Query that matches the similarity memory prior $\left\{{E}_{i}={E}_{i1},{E}_{i2},\cdots,{E}_{ic}\mid{E}_{i}^{T}\in{\mathbb{R}}^{C\times M}\right\}$, then ${E}_{i}$ will perform self-attention with ${X}^{i}$, as shown in Equation 3.
${A}_{i}=SM({X}^{i}\cdot{E}_{i}^{T})\cdot{E}_{i}$ (3)
Where $i=\{1,2,3,...\}$, indicating the number of segmented categories, and $SM$ represents the softmax activation function. This step simulates the process of the “grandmother cell” recognizing specific visual stimuli, so the ${E}_{i}$ plays both the role of Key and Value. Then, we concatenate ${A}_{i}$ to obtain semantic category features $A$. Finally, we aggregate semantic category features and global features, and then map to the original feature space through the convolution operation $\psi(\cdotp)$, as shown in Equation 4.
${X}_{2}=\psi(A+{X}^{\prime})$ (4)
After the initialization of the memory bank is completed, it still needs to continuously update it to ensure that it can adapt to complex medical images. At present, the common method is to use a random update strategy, that is, randomly insert a portion of the input features into the corresponding semantic clusters in the memory. This method is not suitable for complex and ever-changing medical images and may introduce interference information. Therefore, we propose a Weight-Loss Dynamic (W-LD) update strategy to dynamically update the memory bank, as shown in Figure 4(b). Specifically, assume that ${X}^{i}\in{\mathbb{R}}^{N\times C}$ in the output ${X}_{2}\in{\mathbb{R}}^{H\times W\times C}$ of the DMW-LA matches the similarity memory prior ${E}_{i}\in{\mathbb{R}}^{M\times C}$. Then use the Softmax to calculate the weights of ${X}^{i}$ and ${E}_{i}$ separately, and sort the columns of ${X}^{i}$ and ${E}_{i}$ respectively according to the weight size. Note that the columns with large weight in ${X}^{i}$ contain more semantic information, while the columns with small weight in ${E}_{i}$ contain less semantic information. Therefore, we replace the $K$ columns in ${E}_{i}$ with the smallest weight with the $K$ columns in ${X}^{i}$ with the largest weight. In addition, we evaluate the quality of the current prototype memory bank by calculating the difference between the current epoch loss (${Loss}^{\varDelta}$) and the previous epoch loss (${Loss}^{\nabla}$). Specifically, when the value of ${Loss}^{\nabla}-{Loss}^{\varDelta}$ is greater than 0, it shows that the prototype memory bank has a positive impact on network training, leading to the reduction of the current epoch loss. Therefore, the $K$ should be reduced to preserve these valuable features. Conversely, when the value of ${Loss}^{\nabla}-{Loss}^{\varDelta}$ is less than 0, the $K$ should be increased to expand the range of updates and eliminate useless features as much as possible. In this way, the prototype memory bank can accumulate high-value semantic information to the maximum extent in each training epoch. In addition, a linear function is designed to link $K$ and ${Loss}^{\nabla}-{Loss}^{\varDelta}$, as shown in Equation 5.
Figure 5: Overview of DS-GIM. DS-GIM uses cosine similarity and euclidean distance to comprehensively evaluate and highlight the differences between spatial element structures.
$K=(-\alpha\cdot({Loss}^{\nabla}-{Loss}^{\varDelta})+\beta)\cdot\theta$ (5)
In the above formula, $\alpha$ and $\beta$ are decay factors, which are set to 0.5, and $\theta$ = $M$. In addition, we set $\frac{M}{4}\leq K\leq\frac{3M}{4}$ to avoid $K$ falling into extreme situations and ensure stable training, see appendix A. In this way, we can not only minimize the loss of information but also ensure the semantic clarity and representation ability of the similarity memory prior.
3.3 Double-Similarity Global Internal Enhancement Module
The structure of the Double Similarity Global Internal Enhancement Module (DS-GIM) is shown in Figure 5. Assuming the input size is $X\in{\mathbb{R}}^{H\times W\times C}$, we first divide $X$ into $\left\{{X}_{1},{X}_{2},\cdots,{X}_{\frac{H\times W}{{p}^{2}}}\mid{X}_{i}\in{\mathbb{R}}^{P\times P\times C}\right\}$ with the windows of $P\times P$, and calculate the internal cosine similarity of these windows. Weight different decay coefficients for each window based on similarity (from large to small). The purpose is to use the decay coefficient to increase the difference between windows. The formula for calculating the decay coefficient is as follows.
$\gamma=\exp\left(-\left(0.25-2^{-2.5-\frac{5\cdot\mathbf{n}}{\text{$l$}}}\right)\right)$ (6)
In the above formula, $l$ represents the total number of layers in the current stage, $n=\{0,1,2,\cdots,\text{$l$}-1\}$. Next, concatenate these windows to form a new feature representation ${X}^{{}^{\prime}}$. Map ${X}^{{}^{\prime}}$ to a new feature space for feature calculation to obtain ${X}_{I}\in{\mathbb{R}}^{HW\times HW}$. Then average the columns of ${X}_{I}$ to get ${X}_{II}\in{\mathbb{R}}^{HW\times 1}$, and then copy ${X}_{II}$ to get a new matrix $S\in{\mathbb{R}}^{HW\times HW}$. At this time, the elements in $S$ can be regarded as a set of spatial location elements, and we can calculate the euclidean distance between elements using the formula $Euc=\left|S-{S}^{T}\right|$ to measure their proximity in the feature space. Based on size, we can divide the internal elements of $Euc$ into ${Ecu}_{I}$ and ${Ecu}_{II}$. Next, these two parts are calculated by the Softmax function and are hadamard products with ${X}_{I}$ respectively. Finally, splice these two parts together to form a new matrix and multiply it with ${X}^{{}^{\prime}}$, as shown in Equation 7.
${O}_{D}={X}^{\prime}\cdot(SM({Ecu}_{I})*{X}_{I}+SM({Ecu}_{II})*{X}_{I})$ (7)
In the above equation, * represents hadamard product, and $SM$ represents softmax activation function. In this way, distance information between spatial elements will be embedded into feature encoding to enhance the model’s understanding of the internal structure of input data.
Type Methods ACDC SegPC-2021 ISIC-2018
DSC↑ HD95↓ RV Myo LV DSC↑ HD95↓ Cytoplasm Nucleus DSC↑ HD95↓
CNN U-Net [30] 90.14 1.48 89.61 86.58 94.23 80.54 35.89 80.77 80.30 89.49 4.66
Att-UNet [25] 90.48 1.24 89.15 87.29 95.01 80.40 36.04 79.56 81.37 89.58 5.11
U-Net++ [48] 89.54 1.25 87.97 86.62 94.05 80.07 33.18 79.40 80.12 89.19 4.88
EMCAD [28] 91.43 1.13 90.65 88.78 94.86 80.38 34.84 79.49 81.27 90.03 4.33
MADGNet [24] 91.13 1.17 90.01 88.30 95.07 80.02 33.21 79.55 80.48 89.77 4.30
SelfReg-UNet [49] 91.43 1.19 88.92 89.49 95.88 77.60 35.65 77.44 77.75 88.99 5.74
DconnNet [45] 89.96 1.80 89.16 86.12 94.58 79.84 33.07 79.00 80.68 90.01 4.10
ViT SwinUNet [6] 89.49 2.03 88.01 86.03 94.45 78.22 35.24 77.33 79.11 87.32 5.47
TransDeepLab [1] 89.22 2.07 87.48 85.74 94.44 78.81 33.02 77.38 80.24 90.11 4.07
DAE-former [2] 89.78 1.91 89.91 84.38 95.04 78.52 34.86 77.10 79.93 89.85 4.22
SelfReg-SwinUNet [49] 91.49 1.97 89.49 89.27 95.70 76.40 32.45 74.70 78.09 89.16 4.90
CNN& ViT TransUNet [8] 89.37 1.48 89.20 85.23 93.67 76.69 33.95 78.69 80.68 89.92 4.34
MixTrans [38] 90.98 2.28 89.52 88.60 94.82 80.72 32.62 79.78 81.66 86.46 7.34
TransClawUNet [46] 91.29 1.10 91.31 87.93 94.64 79.77 34.45 79.04 80.50 89.62 4.81
HiFormer [18] 90.12 2.15 91.06 84.54 94.77 79.54 32.08 77.73 81.35 89.79 4.58
TransCASCADE [27] 91.48 1.12 91.37 88.32 94.76 80.36 32.87 79.18 81.53 89.76 4.42
PAG-TransYnet [5] 91.33 1.15 90.80 88.28 94.92 79.41 33.88 78.67 80.16 89.65 4.53
Memory Mechanism Sim-MPNet (Ours) 92.18 1.08 92.06 89.19 95.29 83.12 29.66 82.00 84.24 90.85 3.55
Table 1: Performance on ACDC, SegPC-2021 and ISIC-2018. We report the DSC and HD95 scores of different networks on the ACDC, SegPC-2021 and ISIC-2018. We adopt publicly available SOTA methods for implementation, and all results are obtained under the same dataset partitioning conditions. DSC are reported for individual organs or lesions. ↑(↓) denotes the higher (lower) the better.
Type Methods P(M) F(G) DSC↑ HD95↓ Aorta GB KL KR Liver PC SP SM ${T}_{i}$
CNN U-Net [30] 31.04 50.14 76.85 39.70 89.07 69.72 77.77 68.60 93.43 53.98 86.67 75.58 729
Att-UNet [25] 34.88 34.88 77.77 36.02 89.55 68.88 77.98 71.11 93.57 58.04 87.30 75.75 761
U-Net++ [48] 36.63 107.01 78.98 30.65 87.45 65.86 81.75 74.27 93.85 64.16 87.82 76.65 1139
EMCAD [28] 26.76 4.29 83.63 15.68 88.18 68.87 88.08 84.10 95.26 68.51 92.17 83.92 752
MADGNet [24] 32.16 7.71 80.98 27.46 88.01 65.12 82.81 79.09 94.92 67.23 90.15 80.50 1152
SelfReg-UNet [49] 17.26 30.71 80.34 26.98 88.74 71.78 85.32 80.71 93.80 62.21 84.78 75.39 625
DconnNet [45] 36.40 4.80 81.56 19.01 87.48 68.56 83.70 81.04 94.64 65.24 90.64 81.15 849
ViT SwinUNet [6] 27.17 5.92 79.13 21.55 85.47 66.53 83.28 79.61 94.29 56.58 90.66 76.60 625
TransDeepLab [1] 28.61 17.08 80.16 21.25 86.04 69.16 84.08 79.88 93.53 61.19 89.00 78.40 830
DAE-Former [2] 48.07 26.07 82.43 17.46 88.96 72.30 86.08 80.88 94.98 65.12 91.94 79.19 1187
SelfReg-SwinUNet [49] 27.15 5.92 80.54 20.10 86.07 69.65 85.12 82.58 94.18 61.09 87.42 78.22 1014
CNN& ViT TransUNet [8] 105.32 24.66 77.48 31.69 87.23 63.13 81.87 77.02 94.08 55.86 85.08 75.62 794
MixTrans [38] 79.07 44.75 78.59 26.59 87.92 64.99 81.47 77.29 93.06 59.46 87.75 76.81 958
TransClawUNet [46] 113.02 38.08 78.09 26.38 85.87 61.38 84.83 79.36 94.28 57.65 87.74 73.55 1253
HiFormer [18] 29.52 11.88 80.69 19.14 87.03 68.61 84.23 78.37 94.07 60.77 90.44 82.03 931
TransCASCADE [27] 123.47 56.93 82.68 17.34 86.63 68.48 87.66 84.56 94.43 65.33 90.79 83.52 1472
PAG-TransYnet [5] 144.22 33.64 83.43 15.82 89.67 68.89 86.74 84.88 95.87 68.75 92.01 80.66 940
Memory Mechanism Sim-MPNet (Ours) 91.30 43.74 84.34 14.85 87.06 73.49 88.65 85.34 95.28 69.25 91.36 84.30 873
Table 2: Performance on Synapse. P and F represent Parameters and FLOPs respectively. FLOPs of all the methods are reported for 224$\times$224 inputs, except Sim-MPNet (dual encoder: 224$\times$224, 256$\times$256). ${T}_{i}$ is the average inference efficiency of a single image in milliseconds. DSC are reported for individual organs. GB, KL, KR, PC, SP and SM represent gallbladder, left kidney, right kidney, pancreas, spleen and stomach respectively.
4 Experiments
4.1 Datasets
We conduct extensive experiments on four publicly available datasets: Automated Cardiac Diagnosis Challenge (ACDC, 4 classes) [4], Segmentation of Multiple Myeloma Plasma Cells in Microscopic Images (SegPC-2021, 3 classes) [17, 13, 16], International Skin Imaging Collaboration 2018 Challenge (ISIC-2018, 2 classes) [10], and MICCAI 2015 Multi-Atlas Abdomen Labeling Challenge (Synapse, 9 classes). (1) ACDC dataset consists of Magnetic Resonance Imaging (MRI) of the left ventricle (LV), right ventricle (RV) and myocardium (Myo). The division of training set and testing set is consistent with TransUNet++ [42]. (2) SegPC-2021 dataset consists of myeloma plasma cell images. The segmentation target is the cytoplasm and nucleus of the cells. We divide the training set and the testing set by 4:1 ratio. (3) ISIC-2018 dataset consists of dermoscopy images. We also divide the training set and the testing set by 4:1 ratio. (4) Synapse dataset consists of axial abdominal clinical CT images. Each image contains part or all of the eight abdominal organs (aorta, gallbladder, left kidney, right kidney, liver, pancreas, spleen, stomach). We divide the training set and the testing set according to the TransUNet [8].
Figure 6: Multi-organ and cell segmentation. Comparison of visualization results between Sim-MPNet and other SOTA methods on Synapse and SegPC-2021. The different segmentation targets are represented by different colors.
4.2 Implementation Details and Evaluation Metrics
In all experiments, the hyperparameters of DS-GIM (decay coefficient) and DMW-LA (linear function) remained unchanged. For different datasets, just set k in K-means as the number of categories in the dataset. The input resolutions of two encoders are set to 224$\times$224 and 256$\times$256 respectively. In addition, we train the model using an AdamW optimizer with a learning rate of 0.0001, and a weight decay of 1e-4. The model is trained on a single RTX 3090 GPU with a batch size 20 using Dice ($L_{dice}$) and Cross-entropy ($L_{ce}$) losses for 300 epochs. The specific proportion of these two losses is $Loss=0.7\cdot L_{dice}+0.3\cdot L_{ce}$. Finally, all experiments are performed in the Python 3.10.0 and Torch 2.3.0 environment. Only use random rotation and flipping as data augmentation. We choose the Dice score (DSC) and 95% Hausdorff distance (HD95) as evaluation metrics. The HD95 represents the ${95}^{th}$ percentile distance between the boundaries of the model prediction graph and real labels.
4.3 Comparison with SOTA Methods
ACDC. ACDC datasets test the model’s ability to explore local details, and ViT-based methods are often difficult to achieve better segmentation results. Table 1 shows that the DSC of SwinUNet, TransDeepLab and DAE-Former does not exceed 90%. Although the hybrid CNN-ViT methods partially compensate for the shortcomings of ViT, its highest DSC is only 91.48%. In contrast, Sim-MPNet can directly capture the category features of organs in ACDC by dynamic memory prior mechanism, thus achieving 92.18% DSC and acquiring better segmentation results.
SegPC-2021. Table 1 shows that the Sim-MPNet outperforms existing methods and achieves a new SOTA on SegPC-2021. Because most of the myeloma plasma cells in SegPC-2021 are closely connected, this requires the network to have a strong ability to extract local details. Therefore, we can observe that the CNN-based methods and the hybrid CNN-ViT methods are better than the ViT-based methods in segmentation effect. However, compared with the optimal network MixTrans in the above methods, Sim-MPNet improved DSC by 2.40% and HD95 by 2.42mm. This proves that Sim-MPNet can effectively handle complex medical image.
ISIC-2018. Table 1 shows that the Sim-MPNet also achieved SOTA results on the ISIC-2018 dataset. Since there are many large-scale lesions in the ISIC-2018, it verifies the global modeling ability of the network. Thus, the ViT-based TransDeepLab method has achieved the best segmentation results among the existing methods. However, compared with TransDeepLab method, Sim-MPNet improves DSC by 0.74% and HD95 by 0.52mm. This proves once again that Sim-MPNet has the advantage of capturing category dependencies.
Synapse. To further verify the performance of Sim-MPNet model in different complex scenarios, we introduce more diverse Synapse datasets for evaluation. Table 2 shows the comparison results on the Synapse. Firstly, CNN-based methods have significant advantages in small organ segmentation due to their excellent local modeling ability. For example, EMCAD achieved DSC of 88.08% and 84.10% on the left kidney (Kidney (L)) and right kidney (Kidney (R)), respectively. In contrast, Sim-MPNet performed better on these two organs, with DSC of 0.57% and 1.24% higher than EMCAD, respectively. However, on the Spleen with a complex contour, Sim-MPNet cannot fully calculate the correlation between local elements, and thus it cannot achieve the optimal segmentation result. Secondly, ViT-based methods have advantages in large-scale organ segmentation due to their global context exploration ability. DAE-Former achieved DSC of 72.30% and 94.98% on Gallbladder and Liver, respectively. Sim-MPNet has a DSC of 1.19% and 0.3% higher than DAE-Former in these two organs, respectively. Finally, the hybrid CNN-ViT methods combine the advantages of CNN and ViT, enabling the PAG-TransYnet network to explore both local and global features simultaneously, achieving optimal segmentation results in Aorta with the smallest area and Liver with the largest area. Although Sim-MPNet is slightly inferior to PAG-TransYnet in segmentation of the Aorta and Liver, its unique dynamic memory prior mechanism enables it to achieve the best average DSC and HD95 scores on the Synapse. This indicates that Sim-MPNet has excellent robustness in various scenarios. In addition, Sim-MPNet does not have significant advantages in terms of Params and FLOPs, but our method extracts medical image features from a new perspective and achieves excellent segmentation results. This shows the application potential of our method in the field of medical image segmentation and opens up a new path for future research.
Visual Comparison and Analysis. Figure 6 shows the visualizations of Sim-MPNet and other networks on Synapse and SegPC-2021. Firstly, the segmentation results of Sim-MPNet are very close to the ground truth in Synapse. Especially when dealing with organs like the Pancreas, which have complex contours and rich details, Sim-MPNet shows better segmentation performance compared to other networks. Secondly, Sim-MPNet still obtains the best segmentation results for images containing multiple cells in SegPC-2021. Although false positives and false negatives may also occur, the incidence of these errors is significantly reduced compared to other networks. In addition, the semantic clustering receptive fields of DMW-LA on Synapse, ACDC and SegPC-2021 see appendix B.
Encoder 1 Encoder 2 ACDC SegPC-2021 ISIC-2018 Synapse
DMW-LA (Ours) 91.54 82.65 89.57 83.46
MaxViT 91.06 81.92 89.23 83.27
MaxViT MaxViT 91.33 82.07 90.04 83.59
DMW-LA (Ours) MaxViT 92.18 83.12 90.85 84.34
Table 3: Ablation study of DMW-LA on four datasets. We use different encoders to verify their effectiveness separately. The input resolution of each encoder is the same as that of the comparative experiment, which is 224$\times$224 (Encoder 1) and 256$\times$256 (Encoder 2) respectively.
4.4 Ablation Study
Effectiveness of DMW-LA. In this paper, Sim-MPNet adopts dual encoder structure, where Encoder 1 is composed of DMW-LA and Encoder 2 is composed of MaxViT. To evaluate the effectiveness of DMW-LA, we conducted experiments on each encoder. As shown in Table 3, the obvious performance difference indicates that DMW-LA can help the network obtain rich category features, thereby improving segmentation performance.
Effectiveness of DS-GIM and W-LD Update Strategy. We conducted ablation experiments on DS-GIM and W-LD update strategy. It can be seen from Table 4 that compared with the random update strategy, our W-LD update strategy can minimize redundant information and retain useful features. In addition, the remarkable performance improvement in the experiment further confirms the effectiveness of DS-GIM in capturing the internal differences of features.
Dataset ACDC SegPC-2021 ISIC-2018 Synapse
w/o W-LD 91.53 82.63 90.22 83.75
w/o DS-GIM 91.67 82.58 90.47 83.64
w/o W-LD&DS-GIM 91.02 82.23 90.07 83.39
Sim-MPNet 92.18 83.12 90.85 84.34
Table 4: Ablation study of DS-GIM and W-LD update strategy on four datasets. In w/o W-LD, we use a random update strategy to replace the W-LD update strategy. The figures in the table are DSC.
5 Conclusion
In this paper, we break the limitations of traditional methods and construct a novel medical image segmentation network, namely Sim-MPNet. Sim-MPNet directly extracts and locates the category features of lesions or organs. We develop two core modules: DMW-LA and DS-GIM. DMW-LA obtains category information through prototype memory bank and W-LD update strategy, while DS-GIM enhances the intrinsic differences between elements through a dual similarity mechanism. The comparative experiments show that Sim-MPNet outperforms the previous SOTA methods in segmenting multiple organs, cells, and lesions. We believe that Sim-MPNet brings new perspectives and inspirations to the design of medical image segmentation networks in future.
Acknowledgements
This work was supported in part by the Tianshan Talent Training Program under Grant 2022TSYCLJ0036, in part by the National Natural Science Foundation of China under Grant 62472368, Grant 62302427 and Grant 62462060.
Supplementary Material
Appendix A Setting about $K$ in W-LD Update Strategy
To accurately locate the category features of the input image, we propose the W-LD update strategy to dynamically update the prototype memory bank. The following is an introduction to the $K$ parameter in the W-LD update strategy.
Figure 7: The variation curve of $K$ on the Synapse dataset. $M$ is the feature dimension of similarity memory prior.
In W-LD update strategy, $K$ represents the update range of prototype memory bank. To prevent model training collapse, we add a restriction to $K$ ($\frac{M}{4}\leq K\leq\frac{3M}{4}$) to ensure effective updates to the prototype memory bank. It has been proven that our measure is effective, and there is no training collapse in all comparative experiments and ablation experiments. Figure 7 show the curve of $K$ in the training process. It can be seen that with the increase of epoch, the model gradually converges, and $K$ is stable around $\frac{M}{2}$. In the initial stage of training, $K$ fluctuates violently, indicating that the model is learning useful patterns in data, and both large and small initial $K$ may have negative effects. Thus, we set the initial value of $K$ to $\frac{M}{2}$ in all experiments to cooperate with model training.
Appendix B Additional Visualization Results
In the process of medical image segmentation, we assign a semantic cluster (similarity memory prior) to the segmentation target in each image to extract key categorical features. Figure 8 shows the semantic cluster receptive field of DMW-LA on three datasets. We can observe that different clusters encode different semantic concepts. In the Synapse dataset, Cluster 2, Cluster 4 and Cluster 6 tend to focus on the region where the organs are located. In the ACDC dataset, Cluster 1 tends to focus on the left ventricle, while Cluster 2 focuses more on the right ventricle. In the SegPC-2021 dataset, Cluster 1 tends to focus on cytoplasm, while Cluster 3 focuses more on tissue fluid and other cells in the background. This distribution indicates that when processing images, each semantic cluster focuses on specific category features, thereby helping the model identify key structures in medical images.
Figure 8: Visualization of semantic clusters (similarity memory priors) using DMW-LA on Synapse, ACDC, and SegPC-2021. We calculate the distribution of similarity memory prior in the prototype memory bank, and highlight their receptive fields with white boxes.
Appendix C Evaluation Metrics
To better evaluate the performance of the network, we choose Dice score (DSC) and 95% Hausdorff distance (HD95) as evaluation metrics, as shown in Equation 8 and Equation 9.
${DSC=\frac{(2*TP)}{(2*TP+FP+FN)}}$ (8)
[…]
$\max_{b\in B}\{\min_{a\in A}\{d(b,a)\}\}\}$
In the above formula, $TP$ represents foreground information is predicted as foreground, $FP$ represents background information is predicted as foreground, and $FN$ represents foreground information is predicted as background. The HD95 represents the ${95}^{th}$ percentile of distances between the boundaries of A and B. DSC is used to evaluate the degree of overlap, while HD95 is used to evaluate distance. The combination of them can comprehensively evaluate the performance of the model.